The solution to this system of equations is (-2,1). Find the values of p and q
px + (9-q)y=-10
(3p+1)-(q-6)y=-7

Answer by ikleyn(52790)

.
The solution to this system of equations is (-2,1). Find the values of p and q
px + (9-q)y=-10
(3p+1)-(q-6)y=-7
--------------------------------------

Substitute the given solution x = -2 and y = 1 into the system. You will get

(-2)*p + 1*(9-q) = -10,
3p + 1 - 1*(q-6) =  -7.

Simplify:

-2p - q = -10 - 9,
3p - q  = -7 - 1 - 6,

or

-2p - q = -19.
 3p - q = -14.

It is your system for p and q.

Now solve it to find p and q. For it, distract the first equation from the second one.

The next steps are obvious. Please complete the solution yourself.